How many gallons of a 60% antifreeze solution must be mixed with 70 gallons of 25% antifreeze to get a mixture that is 50% antifreeze? Use 6 step method.
Answer by mananth(16946)

----------- percent ---------------- quantity
Antifreeeze I 25 ---------------- 70 L
Antifreeze II 60 ---------------- x L
Total 50 ---------------- 70 + x L

25*70+60*x=50(70+x)

1750+60x=3500+50x

60x-50x=3500-1750

10x=1750

/10

x=175 L 60 % Antifreeze II